I have an exam tomorrow and I forget how to solve a triple matrix with the variables in the middle. The following is an example:
"In the school cafeteria, Larry purchased a cookie, two chicken fingers and a can of Pepsi for $4.30. Curly purchased three cookies, one chicken finger and two cans of Pepsi for $5.30. Moe purchased two cookies, one chicken finger, and two cans of Pepsi for $4.55. What is the individual price of a cookie, a chicken finger, and a can of pop?"

Answer by richwmiller(17219)

I will set it up for you.
1c+2f+1p=430
3c+1f+2p=530
2c+1f+2p=455
1 2 1 430
3 1 2 530
2 1 2 455

cookie c = 75, chicken fingers f = 135, pepsi p = 85